kara and jennifer are selling boxes of chocolate to raise money for their senior trip to the the grand canyon. customers can buy small boxes of chocolates and large boxes of chocolate. kara sold 11 small boxes and 11 large boxes for a total of 220. jennifer sold 3 small boxes and 14 large boxes for a total of 203. what would be my total if i purchase 2 small and 2 large boxes from jennifeer
Answer by mananth(16949)

let cost of small box be $x
large box be $y
KARA
11x+11y=220
/11
x+y =20.................1
Jennifer
3x+14y = 203............2
...
multiply equation 1 by -3
we get
-3x-3y=-60
add this to equation 2
11y = 203-63
11y=143
/11
y=13 price of large box
..
plug value of y in 1
x+y=20
x+13=20
x=20-13
x= 7 price of small box.
..
Price of 2 small & 2 large =
2*7+2*13= $40
..
m.ananth@hotmail.ca